OpenStudy (anonymous):

find the indefinite integral [(1/x^11)-x^11-1/9]dx

OpenStudy (anonymous):

\[\int\limits_{}^{}(1/x ^{11}-x ^{11}-1/9\]

OpenStudy (lgbasallote):

distribute the integral symbol first

OpenStudy (lgbasallote):

then use power rule \(\LARGE x^n = \frac{x^{n+1}}{n+1}\)